Use transformations to graph each of the following y=_(x+5)^2-4


1
Expert's answer
2022-01-27T13:50:07-0500

1. Given the graph


"y_1=x^2"

2. Shift the graph of "y_1=x^2" to the left "5" units


"y_2=(x+5)^2"

3. Shift the graph of "y_2=(x+5)^2" downward "4" units


"y=(x+5)^2-4"